# Worst-Case to Expander-Case Reductions: Derandomized and Generalized

### v1

@@ 2 Technical Overview
In this section, we summarize the core gadget of Abboud and Wallheimer [3] (henceforth, AW) that is used in all their Direct-WTERs, and then present our modification. The techniques in AW’s work are based on a simple core gadget that takes a graph $G$ as input and augments it with vertices and edges to output an expander $G_{exp}$ containing $G$ as an induced subgraph. Roughly, their construction boils down to the following procedure.
@@
Given a graph $G=(V,E)$, add a set $U$ of $n$ vertices called expansion layer. Then, for every vertex $v\in V$, take a sample of $O(\log n)+\deg_{G}(v)$ vertices in $U$ and make them neighbors of $v$. Clearly, the size of this graph is $O(n\log n+m)$. In addition, they prove that it is an $\Omega(1)$-expander with high probability by showing that the probability that a cut in $G_{exp}$ ends up being sparse is very low. Hence, the result follows the union bound. [^10] We remark that their construction also contains some additional $O(n\log n)$ edges, which are added deterministically, but careful inspection shows that they are not needed. The time complexity to construct the expander is $\tilde{O}(n+m)$ time. However, there is no guarantee that the problem’s solution to $G$ can be computed easily from the solutions to $G_{exp}$. To this end, AW provide additional, mostly simple gadgets to control the solution in $G_{exp}$ in a predictable manner while preserving the conductance of $G_{exp}$ up to a constant factor.
@@
The first limitation of AW’s approach that we aim to resolve in this paper is their use of randomness. To do so, one could attempt to introduce an explicit, $d$-regular bipartite $\Omega(1)$-expander $X$ between $V$ and $U$. However, which degree $d$ should we choose? For starters, we consider picking some constant $d\geq 3$. However, this approach might fail if $G$ contains cuts with many internal edges and few outgoing edges. For instance, if $G$ includes a cut $S$ of $\sqrt{n}$ vertices that form an isolated $\sqrt{n}$-clique, then in $G_{exp}$ we have $vol(S)=\Theta(n)$, while the number of outgoing edges added by the expander is $O(|S|d)=O(\sqrt{n})$, thus $\phi(S)=O(1/\sqrt{n})$. On the opposite end, we could select a complete bipartite graph with $\Theta(n^{2})$ edges. While this will address the problem and indeed result in an $\Omega(1)$-expander, this approach is undesirable due to the excessive number of added edges. Hence, the most natural strategy to consider is incorporating a $d$-regular expander for $d=\lceil\frac{m}{n}\rceil$. To be more precise, since explicit constructions exist only for $d\geq 3$, the approach is to pick $d=\lceil\frac{m}{n}\rceil+3$ (as it could be the case that, for example, $m=0$). While the number of added edges is now $O(m+n)$, this approach is slightly naïve, as it does not even address the above problem of isolated $\sqrt{n}$-clique in a graph with $O(n)$ edges. The crux of the issue lies in the fact that $G$ is not regular. If $G$ were $d$-regular, then the inclusion of a $d$-regular expander would resolve our problem. One can look for expanders with a certain degree sequence, matching the one of $G$, but constructing those seems challenging.
@@
The starting point for our modification is the observation that the randomness in AW’s constructions is coming to accomplish two things at once: first and foremost, choosing random edges breaks structure and creates a “random like” graph. However, second, and less obviously, it is a way to achieve a certain “balanced allocation” of neighbors in $U$, resulting in nearly uniform degrees in $U$, which is crucial for their analysis to work. This leads to our modification, in which we substitute their single expansion layer $U$ with two layers: one for balanced allocation denoted $L$, connected to $V$ using a load-balancing algorithm, ensuring that the degrees in this layer are about $\frac{2m}{n}$, and another layer for expansion denoted $R$, connected to $L$ using the edges of a bipartite, $(\frac{2m}{n}+3)$-regular $\Omega(1)$-expander. A random graph can accomplish the construction of the first layer, as AW did, but it could also be accomplished in any other way, such as the standard Round-Robin algorithm without any randomness. Then, all we have to do is use an explicit construction (described in Appendix C) of a regular $\Omega(1)$-expander. This is the whole core gadget in Section 4. Using it, we follow the steps of AW and add various gadgets to $G_{exp}$, such that the solution for $G$ can be retrieved from the solution for $G_{exp}$.
@@
The second limitation of AW’s approach, which we aim to resolve, is that it focuses solely on the static model of computation. To dynamize our deterministic core gadget, we employ a lazy-greedy approach as follows. As edges are inserted and deleted from $G$, we lazily wait until some local or global imbalance event occurs. The local imbalance event is when the degree of a vertex in $G$ doubles itself. In this case, we add more edges from this vertex to $L$ using a (deterministic) greedy balanced allocation strategy. We prove that this greedy strategy ensures that the degrees in $L$ remain within a constant factor from each other for a long time, which is a sufficient condition for the graph to remain an $\Omega(1)$-expander. The global imbalance events are as follows. One event is when the degrees in $L$ become overly imbalanced, in which case we apply a costly rebalance procedure using Round-Robin. Another global imbalance event is when the number of edges in $G$ differs too much from the number of edges in $G_{exp}$, in which case we apply a costly reconstruction procedure using the static core gadget. This approach results in an amortized slowdown of $O(1)$ time per operation.
@@ 3 Preliminaries
Let $G=(V,E)$ be an undirected simple graph. Throughout the paper, we use $n$ and $m$ to denote the number of edges in vertices in $G$, respectively. Denote th neighborhood of $v$ by $N(v):=\{u\in V\mid(v,u)\in E\}$ and the degree of $v$ by $\deg(v)$. We say that $G$ is $d$-regular if $\deg(v)=d$ for all $v\in V$. A vertex $v$ is called a pendant vertex if $\deg(v)=1$. The set of all edges with one endpoint in $S\subseteq V$ and another endpoint in $T\subseteq V$ is denoted by $E(S,T):=\{(u,v)\in E\mid u\in S,v\in T\}$, and its cardinality is denoted by $e(S,T):=|E(S,T)|$. We call $E(S,V\setminus S)$ the out-going edges of $S$. We employ subscripts to indicate which graphs we refer to when it is not clear from the context. For instance, $\deg_{H}(v)$ denotes the degree of vertex $v$ in a graph $H$.
Conductance.
Let $S\subseteq V$ be a cut. The volume of $S$ is defined as $vol(S):=\sum_{v\in S}\deg(v)$. The conductance of $S$ is defined as $\phi(S):=e(S,V\setminus S)/\min(vol(S),vol(V\setminus S))$. If $\min(vol(S),vol(V\setminus S))=0$, we define $\phi(S)=0$. The conductance of the entire graph $G$ is defined as $\phi_{G}:=\min_{\emptyset\neq S\subseteq V}\phi(S)$. Throughout this paper, unless explicitly indicated otherwise, we adhere to the following definition of expander graphs, in which $0\leq\phi\leq 1$.
@@
This definition aligns with the one commonly used in the literature about expander decompositions.
Edge-expansion.
Another notion of expansion that we use in our proofs is edge expansion.
@@ 4.1 The core gadget
    (Degree balancing) For every $v\in V$ add, using a Round-Robin algorithm, $\deg_{G}(v)+3$ neighbors in $L$. Namely, follow a circular order over $L$ to pick neighbors one at a time. Assuming $N\geq n+2$, every vertex is connected to $\deg_{G}(v)+3$ distinct neighbors in $L$ without wrapping around. The Round-Robin algorithm guarantees that the degrees within $L$ are almost balanced. Namely, since the total number of edges that we add in this step is $\sum_{v\in V}(\deg_{G}(v)+3)=2m+3n$, the degree of every vertex in $L$ is either $\lfloor\frac{2m+3n}{N}\rfloor$ or $\lceil\frac{2m+3n}{N}\rceil$.
@@
Denote the resulting graph by $G_{exp}:=(V_{exp},E_{exp})$, illustrated in Figure 1. Note that the running time of this gadget is $\tilde{O}(m+n+dN)=\tilde{O}(m+n)$.
@@
For clarity, we henceforth omit the subscript $G_{exp}$ from $e_{G_{exp}}(\cdot,\cdot)$, and keep the subscript $G$ in $e_{G}(\cdot,\cdot)$. Similarly, we do for $E(\cdot,\cdot)$, $\deg(\cdot)$, and $vol(\cdot)$. Let us now prove the main lemma.
@@
On the high level, we split into two cases. In the first case, we deal with cuts whose volume is mostly concentrated in $X$ and show that there must be many outgoing edges because $X$ is an expander, and the degrees in $L$ are bounded. In the second case, we deal with cuts whose volume is mostly concentrated in $G$ and show that there must be many outgoing $V$-to-$L$ edges in this case. See Figure 2. In more detail:
Proof.
@@ 5.2 Fully-dynamic core gadget
The above core gadget is not suitable for the fully dynamic setting, where $G$ undergoes edge insertions and deletions. Let us discuss the difficulties in dynamizing it and our high-level approach. Two issues arise in this setting. The first one is that when the degree of a vertex $v\in V$ increases due to edge insertions, we need to allocate to $v$ additional neighbors from $L$ while preserving balance in the degrees in $L$. One attempt to solve this issue is to pick a minimum-degree vertex in $L$, which can be computed quickly using a priority queue. While this ensures balanced degrees in $L$, there is a subtle issue that the minimum-degree vertex in $L$ might already be a neighbor of $v$; hence we can not add another edge to it. [^11] In the multigraph setting, this would not become an issue, as we can connect two vertices with multiple edges. However, our focus in this paper is on the more restricted family of simple graphs and simple expanders.
Therefore, we slightly modify this heuristic, and instead, our suggested greedy approach is to compute the successor of the minimum-degree vertex repeatedly, until we find a minimum-degree vertex in $L\setminus N(v)$, and then make it a neighbor of $v$. The required computation here is proportional to the degree of $v$, which might not be constant. Hence, we employ lazy updates; we wait until the degree of $v$ doubles itself before computing a new batch of neighbors in $L$ so that the amortized cost will be constant. One might notice, however, that the minimum-degree vertex in $L\setminus N(v)$ is not necessarily a minimum-degree vertex in $L$, so it needs to be clarified that the degrees in $L$ remain balanced. Nonetheless, we manage to show that the degrees in $L$ become overly imbalanced only after $\Omega(m+n)$ edge insertions. In this case, there is enough credit to reconstruct $E(V,L)$ using the Round-Robin algorithm, resulting in a total amortized time $O(1)$ per edge insertion. In general, edge deletions do not require further computation. However, too many edge deletions will make $X$ too dense compared to $G$, resulting in a large blowup. This is the second core issue.
The second issue is to maintain an appropriate degree of $X$. Recall that the degree of $X$ is picked to a value that is essentially the average degree of $G$. Therefore, if $G$ undergoes too many edge insertions, it may result in $X$ being too sparse, and the conductance of $G^{exp}$ becomes small. On the other hand, if $G$ undergoes too many edge deletions, then $X$ becomes too dense, and the blowup becomes large. To resolve this issue, we utilize the lazy updates approach: we recompute $X$ if its degree is too large or too small. This adds additional amortized cost of $O(1)$. Let us now provide the full details of the dynamized core gadget.
Our algorithm makes use of three procedures: (1) $\tt{UPDATE}$, which computes and adds a batch of edges from a vertex in $G$ to $L$, (2) $\tt{BALANCE}$, which rebalances the degrees in $L$ using Round-Robin, and (3) $\tt{RECOMPUTE}$, which recomputes the graph using the static core gadget. In this section, we use $G^{exp}$ to denote a dynamic $\Omega(1)$-expander output by the dynamized core gadget. We use subscripts to indicate the state of a dynamic graph at a certain time, e.g., $G_{t}$ is the dynamic graph $G$ at time $t$. We begin by describing the algorithm, then the three procedures, and lastly, we show the analysis of the algorithm.
@@
In the preprocessing step, given $G_{0}$, apply the static core gadget to construct an $\Omega(1)$-expander $G_{0}^{exp}$. We store the vertices of $L$ sorted according to their degrees, in a data structure supporting updates and successor queries in $O(1)$ time. [^12] Naïvely this would take $\tilde{O}(1)$ time using standard data structures, but it can be optimized to $O(1)$ since the degrees are integers in the range $[1,2N]$, and our updates only increase or decrease the degree of a vertex by $1$. We omit the details for the sake of brevity.
@@
For every insertion of an edge $uv$ to $G$, begin by inserting $uv$ to $G^{exp}$. Denote by $m_{t}$ the number of edges in $G$ at the last time we applied the $\tt{RECOMPUTE}$ procedure (or $m_{0}$ if we did not apply it). If $m\geq 2m_{t}+n$, apply the $\tt{RECOMPUTE}$ procedure to recompute the graph using the static core gadget and finish.
@@
    If $\deg_{G}(v)<2\deg_{L}(v)$, finish. Otherwise, apply the $\tt{UPDATE}$ procedure to $v$ to add a batch of neighbors to $v$, after which we have $\deg_{L}(v)=\deg_{G}(v)+3$.
@@
  • $\tt{UPDATE}$ Let $k:=\deg_{L}(v)$. Compute $k+3$ minimum-degree vertices in $L\setminus N(v)$, denoted $x_{1},x_{2},\ldots,x_{k+3}$, by repeatedly making successor queries to $L$ and skipping vertices which belong to $N(v)$. For every $x_{i}$, insert an edge $vx_{i}$ to $G^{exp}$. Note that now we have $\deg_{L}(v)=\deg_{G}(v)+3$.
  • $\tt{BALANCE}$ Compute, using Round-Robin, a new set of $V$-to-$L$ edges, denoted $R$, and then replace $E(V,L)$ with $R$. Specifically, insert the edges of $R$ in the order given by the Round-Robin algorithm before removing $E(V,L)\setminus R$, to ensure that the degrees do not vary too much in the intermediate graphs.
  • $\tt{RECOMPUTE}$. Apply the static core gadget to compute a set of edges $R$, and $X^{\prime}$, where $R$ is the set of $V$-to-$L$ edges as before, and $X^{\prime}$ is the expander on $L\cup R$. In particular, $X^{\prime}$ is a $d_{X^{\prime}}$-regular, $\Omega(d_{X^{\prime}})$-edge expander for $d_{X^{\prime}}=\lceil\frac{2m+3n}{N}\rceil$. Replace $X$ with $X^{\prime}$ by first inserting the edges of $X^{\prime}$, and then removing the leftover edges of $X$ which do not belong to $X^{\prime}$. Then insert the edges of $R$ and remove $E(V,L)$, as we did above. We think of $X^{\prime}$ as being the “active” expander at every intermediate step of this procedure.
Now, let us analyze the algorithm, starting with the correctness.
@@
The blowup is clearly $O(m+n)$ because at every intermediate step, we add at most $|R|+|e(X^{\prime})|=O(m+n)$ edges. Let us prove that the conductance is $\Omega(1)$. Observe that:

### v2

@@ 2 Technical Overview
In this section, we summarize the core gadget of Abboud and Wallheimer [3] (henceforth, AW) that is used in all their Direct-WTERs, and then present our modification. Roughly, their construction boils down to the following procedure.
@@
Given a graph $G=(V,E)$, add a set $U$ of $n$ vertices called expansion layer. Then, for every vertex $v\in V$, take a sample of $\deg_{G}(v)+O(\log n)$ vertices in $U$ and make them neighbors of $v$. Clearly, the size of this graph and the running time are both $O(m+n\log n)$. In addition, they prove that it is an $\Omega(1)$-expander with high probability by showing that the probability that a cut in $G_{exp}$ ends up being sparse is very low. However, there is no guarantee that the problem’s solution to $G$ can be computed easily from the solutions to $G_{exp}$. To this end, AW provide additional, mostly simple gadgets to control the solution in $G_{exp}$ in a predictable manner while preserving the conductance of $G_{exp}$ up to a constant factor.
@@
The first limitation of AW’s approach that we aim to resolve in this paper is their use of randomness. To do so, one could attempt to introduce an explicit, $d$-regular bipartite $\Omega(1)$-expander $X$ between $V$ and $U$. However, which degree $d$ should we choose? For starters, we consider picking some constant $d\geq 3$. However, this approach might fail if $G$ contains cuts with many internal edges and few out-going edges. For instance, if $G$ includes a cut $S$ of $\sqrt{n}$ vertices that induce an isolated $\sqrt{n}$-clique, then in $G_{exp}$ we have $vol(S)=\Theta(n)$, while the number of out-going edges added by the expander is $O(|S|d)=O(\sqrt{n})$, thus $\phi(S)=O(1/\sqrt{n})$. On the opposite end, we could select a complete bipartite graph with $\Theta(n^{2})$ edges. While this will address the problem and indeed result in an $\Omega(1)$-expander, this approach is undesirable due to the excessive number of added edges. Hence, the most natural strategy to consider is incorporating a $d$-regular expander for $d=\lceil\frac{m}{n}\rceil$. To be more precise, since explicit constructions exist only for $d\geq 3$, the approach is to pick $d=\lceil\frac{m}{n}\rceil+3$ (as it could be the case that, for example, $m=0$). While the number of added edges is now $O(m+n)$, this approach is slightly naïve, as it does not even address the above problem of isolated $\sqrt{n}$-clique in a graph with $O(n)$ edges. The crux of the issue lies in the fact that $G$ is not regular. If $G$ were $d$-regular, then the inclusion of a $d$-regular expander would resolve our problem. One can look for expanders with a certain degree sequence, matching the one of $G$, but constructing those seems challenging.
@@
The starting point for our modification is the observation that the randomness in AW’s constructions is coming to accomplish two things at once: first and foremost, choosing random edges breaks structure and creates a “random like” graph. However, second, and less obviously, it is a way to achieve a certain “balanced allocation” of neighbors in $U$, resulting in nearly uniform degrees in $U$, which is crucial for their analysis to work. This leads to our modification, in which we substitute their single expansion layer $U$ with two layers: one for balanced allocation denoted $L$, connected to $V$ using a load-balancing algorithm, ensuring that the degrees in this layer are about $\frac{2m}{n}$, and another layer for expansion denoted $R$, connected to $L$ using the edges of a bipartite, $(\frac{2m}{n}+3)$-regular $\Omega(1)$-expander. A random graph can accomplish the construction of the first layer, as AW did, but it could also be accomplished in any other way, such as the standard Round-Robin algorithm without any randomness. Then, all we have to do is use an explicit construction (described in Appendix C) of a regular $\Omega(1)$-expander. This is the whole deterministic core gadget in Section 4. Using it, we follow the steps of AW and add various gadgets to $G_{exp}$, such that the solution for $G$ can be retrieved from the solution for $G_{exp}$.
@@
The above core gadget is not suitable for the fully dynamic setting, where $G$ undergoes edge insertions and deletions. The main issue arising is that when the degree of a vertex $v\in V$ increases due to edge insertions, we need to allocate to $v$ additional neighbors from $L$ while preserving approximately balanced degrees in $L$. One attempt to solve this issue is to add an edge from $v$ to a minimum-degree vertex in $L$, which can be computed quickly using a priority queue. While this ensures balanced degrees in $L$, there is a subtle issue that the minimum-degree vertex in $L$ might already be a neighbor of $v$; hence we can not add another edge to it because that would create a parallel edge (which we aim to avoid). Therefore, we slightly modify this heuristic, and instead, our suggested approach is compute the successor of the minimum-degree vertex in $L$ repeatedly, until a minimum-degree vertex in $L\setminus N(v)$ is reached, and then make it a neighbor of $v$. The required computation here is proportional to the degree of $v$, which might not be constant, but using lazy updates this approach results in amortized cost $O(1)$. However, a possible issue arising in this algorithm is that the minimum-degree vertex in $L\setminus N(v)$ is not necessarily a minimum-degree vertex in $L$, so it needs to be clarified that the degrees in $L$ remain balanced, as otherwise the graph might not be an $\Omega(1)$-expander. Nonetheless, we prove that the degrees in $L$ become overly imbalanced only after $\Omega(m+n)$ edge insertions. At this stage, there is enough credit to reconstruct $G_{exp}$ from scratch, resulting in a total amortized cost of $O(1)$ per edge insertion. In general, edge deletions do not require further computation. However, too many edge deletions or insertions will make $X$ too dense or too sparse compared to $G$, resulting either in a large blowup or small conductance. This may happen only after $\Omega(m+n)$ updates, thus it can also be handled in amortized cost $O(1)$ by periodic reconstruction.
@@ 3 Preliminaries
Let $G=(V,E)$ be an undirected simple graph. Throughout the paper, we use $n$ and $m$ to denote the number of edges in vertices in $G$, respectively. Denote the neighborhood of $v$ by $N(v):=\{u\in V\mid uv\in E\}$ and the degree of $v$ by $\deg(v)$. We say that $G$ is $d$-regular if $\deg(v)=d$ for all $v\in V$. A vertex $v$ is called a pendant vertex if $\deg(v)=1$. The set of all edges with one endpoint in $S\subseteq V$ and another endpoint in $T\subseteq V$ is denoted by $E(S,T):=\{uv\in E\mid u\in S,v\in T\}$, and its cardinality is denoted by $e(S,T):=|E(S,T)|$. We call $E(S,V\setminus S)$ the out-going edges of $S$. We employ subscripts to indicate which graphs we refer to when it is not clear from the context. For instance, $\deg_{H}(v)$ denotes the degree of vertex $v$ in a graph $H$.
Conductance and edge-expansion.
Let $S\subseteq V$ be a cut. The volume of $S$ is defined as $vol(S):=\sum_{v\in S}\deg(v)$. The conductance of $S$ is defined as $\phi(S):=e(S,V\setminus S)/\min(vol(S),vol(V\setminus S))$. If $\min(vol(S),vol(V\setminus S))=0$, we define $\phi(S)=0$. The conductance of the entire graph $G$ is defined as $\phi_{G}:=\min_{S\subseteq V}\phi(S)$. Throughout this paper, unless explicitly indicated otherwise, we adhere to the following definition of expander graphs, in which $0\leq\phi\leq 1$.
@@
Another related notion of expansion that we use in our proofs is edge expansion.
@@ 4.1 The core gadget
    (Degree balancing) For every $v\in V$ add, using a Round-Robin algorithm, $\deg_{G}(v)+3$ neighbors in $L$. Namely, follow a circular order over $L$ to pick neighbors one at a time. Assuming $N\geq n+2$, every vertex is connected to $\deg_{G}(v)+3$ distinct neighbors in $L$ without wrapping around. The Round-Robin algorithm guarantees that the degrees within $L$ are almost balanced. In more detail, since the total number of edges that we add in this step is $\sum_{v\in V}(\deg_{G}(v)+3)=2m+3n$, the degree of every vertex in $L$ is either $\lfloor\frac{2m+3n}{N}\rfloor$ or $\lceil\frac{2m+3n}{N}\rceil$.
@@
Denote the resulting graph by $G_{exp}:=(V_{exp},E_{exp})$, illustrated in Fig. 1. Note that the running time of this gadget is $\tilde{O}(m+n+dN)=\tilde{O}(m+n)$.
@@
The proof of this lemma, on the high level, splits into two cases. In the first case, we deal with cuts whose volume is mostly concentrated in $X$, and show that there must be many out-going edges because $X$ is an expander and the degrees in $L$ are bounded. In the second case, we deal with cuts whose volume is mostly concentrated in $G$ and show that there must be many out-going $V$-to-$L$ edges in this case. Let us now prove this claim formally. For clarity, we henceforth omit the subscript $G_{exp}$ from $e_{G_{exp}}(\cdot,\cdot)$, and keep the subscript $G$ in $e_{G}(\cdot,\cdot)$. Similarly, we do for $E(\cdot,\cdot)$, $\deg(\cdot)$, and $vol(\cdot)$.
Proof of Lemma 9.
@@ 5.2 Fully-dynamic core gadget
Our algorithm makes use of three procedures: (1) $\tt{UPDATE}$, which computes and adds a batch of edges from a vertex in $G$ to $L$, (2) $\tt{BALANCE}$, which rebalances the degrees in $L$ using Round-Robin, and (3) $\tt{RECOMPUTE}$, which recomputes the graph using the static core gadget. In this section, we use $G^{exp}$ to denote a dynamic $\Omega(1)$-expander output by the dynamic core gadget. We use subscripts to indicate the state of a dynamic graph at a certain time, e.g., $G_{t}$ is the dynamic graph $G$ at time $t$. Let us describe the algorithm.
@@
In the preprocessing step, given $G_{0}$, apply the static core gadget to construct an $\Omega(1)$-expander $G_{0}^{exp}$. Store the vertices of $L$ sorted according to their degrees, in a data structure supporting updates and successor queries in $O(1)$ time. [^8] Naïvely this would take $\tilde{O}(1)$ time using standard data structures, but it can be optimized to $O(1)$ since the degrees are integers in the range $[1,2N]$, and our updates only increase or decrease the degree of a vertex by $1$.
@@
For every insertion of an edge $uv$ to $G$, begin by inserting $uv$ to $G^{exp}$. Denote by $m_{t}$ the number of edges in $G$ at the last time we applied the $\tt{RECOMPUTE}$ procedure (or $m_{0}$ if we did not apply it yet). If $m\geq 2m_{t}+n$, apply the $\tt{RECOMPUTE}$ procedure to recompute the graph using the static core gadget and finish.
@@
    If $\deg_{G}(v)<2\deg_{L}(v)$, finish. Otherwise, apply the $\tt{UPDATE}$ procedure to $v$ to add a new batch of neighbors of $v$ in $L$, after which we have $\deg_{L}(v)=\deg_{G}(v)+3$.
@@
  • $\tt{UPDATE}$ Let $k:=\deg_{L}(v)$. Compute $k+3$ minimum-degree vertices in $L\setminus N(v)$, denoted $x_{1},x_{2},\ldots,x_{k+3}$, by repeatedly making successor queries to the minimum degree vertex in $L$ and skipping vertices which belong to $N(v)$. For every $x_{i}$, insert an edge $vx_{i}$ to $G^{exp}$. Note that now we have $\deg_{L}(v)=\deg_{G}(v)+3$.
  • $\tt{BALANCE}$ Compute, using Round-Robin, a new set of $V$-to-$L$ edges, denoted $A$, and then replace $E(V,L)$ with $A$. This is done by inserting the edges of $A$ in the order given by the Round-Robin algorithm before removing $E(V,L)\setminus A$, to ensure that the degrees do not vary too much in the intermediate graphs.
  • $\tt{RECOMPUTE}$. Apply the static core gadget to compute a set of edges $A$, and $X^{\prime}$, where $A$ is the set of $V$-to-$L$ edges define above, and $X^{\prime}$ is the expander on $L\cup R$. In particular, $X^{\prime}$ is a $d_{X^{\prime}}$-regular, $\Omega(d_{X^{\prime}})$-edge expander for $d_{X^{\prime}}=\lceil\frac{2m+3n}{N}\rceil$. Replace $X$ with $X^{\prime}$ by first inserting the edges of $X^{\prime}$, and then removing the leftover edges of $X$ which do not belong to $X^{\prime}$. Then insert the edges of $A$ and remove $E(V,L)$, as we did above.
Let us now analyze the dynamic core gadget.
@@
The blowup is clearly $O(m+n)$ because at every intermediate step, we add at most $|A|+|e(X^{\prime})|=O(m+n)$ edges. Let us prove that the conductance is $\Omega(1)$. Observe that: